OpenStudy (anonymous):

Which expression is equivalent to 100n2 – 1? (10n)2 – (1)2 (10n2)2 – (1)2 (50n)2 – (1)2 (50n2)2 – (1)2

OpenStudy (anonymous):

your a bit*ch

OpenStudy (anonymous):

bbbbbbbbbbbbbbiiiiiiiiiittttttttttttttttttccccccccccccccccchhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

bbbbbbbbiiiiiiiittttttccccchhhhhhhhh

jhonyy9 (jhonyy9):

100n^2 -1 = so just us formula a^2 -b^2 =(a-b)(a+b) when a=10n and b=1 hope this will help you